displacement of a particle of mass 2 kg moving in a straight line varies with time as
x=(2t^3 +2)m impulse of the force acting on the particle over a time interval between t=0 and t=1s is

Grade:11

1 Answers

Dinan jana V D
6 years ago
For varying force , impulse I=integral F.dtF=ma=m(d^2x/dt^2) =2(d^2/dt2 of 2t^3+2) =2 (12t) =24tTherefore impulse=integral (24t.dt)=24/2=12.